A ladder is resting against a wall. The top of the ladder touches the wall at a height of 12 feet. Find the length of the ladder if its length is 4 ft. more than its distance from the wall.
Answer by Greenfinch(383)

Distance from wall is X feet. Height up wall is 12 feet.
X^2 + 12^2 = (X + 4)^2
X^2 + 144 = X^2 + 8X + 16
8X + 16 = 144
8X = 128
X= 16
16^2 + 12^2 = 20^2 (3,4,5 triangle)
Length is 20 feet